\begin{thebibliography}{99}
\bibitem{Ru}D. Rudolph.  An example of measure-preserving map with minimal self-joinings, and applications. J. d'Analyse Math., 35 (1979), 97-122. 	
\bibitem{JRS} 	A. del Junco, A.M. Rahe, L. Swanson. Chacon's automorphism has minimal self-joinings. J. d'Analyse Math. 37 (1980), 276-284.	

\bibitem{JR} A. del Junco and D. J. Rudolph.  A rank-one, rigid, simple, prime map.
    Ergod. Th. Dynam. Sys. 7 (1987),  229-247

\bibitem{JR1}del Junco and D.J. Rudolph, On ergodic actions whose self-joinings are graphs, Ergod. Th. Dynam. Sys. 7 (1987), 531-557

\bibitem{K} J. King  Joinings-rank and the structure of finite rank mixing transformation. J. d'Analyse Math., 51 (1988), 182-227
\bibitem{KT}J. L. King, J.-P. Thouvenot. A canonical structure theorem for finite joining-rank maps, J. Analyse Math. 56 (1991), 211-230.
%\bibitem{R92} V.V.  Ryzhikov.  Mixing, rank, and minimal self-joining of %actions with an invariant measure. 
% Sb. Math, 1993, 75:2, 405-427
\bibitem{T}J.-P. Thouvenot. Some properties and applications of joinings in ergodic theory, Ergodic theory and its connections with harmonic analysis, Proc. of the 1993 Alexandria Conference, LMS Lecture notes series, 205, eds. K. Petersen, I. Salama, Cambridge Univ. Press, Cambridge, 1995.
\bibitem{R} V.V.  Ryzhikov. Weak limits of powers, the simple spectrum of symmetric products and mixing
constructions of rank 1, Sb. Math. 198 (2007), 733-754.
\bibitem{FLM}  K. Fraczek, M. Lemanczyk and E. Lesigne.   Mild mixing property for special flows under piecewise constant functions.
Discrete and Continuous Dynamical Systems, 19(2007),  4, 691 - 710. 

\bibitem{B}J. Bourgain, On the correlation of the Moebius function with random rank-one systems, preprint 2011, available at arxiv.org/abs/1112.1032 to appear in J. Analyse

\bibitem{JRR}  E. Janvresse, T. de la Rue and  V.V. Ryzhikov.    Around King's Rank-One Theorems: Flows and $Z^n$-Actions.  Contemporary Mathematics 567 (2012), pp. 143-161

\bibitem{JPRR}  E. Janvresse, A. A. Prikhodko, T. de la Rue
and  V.V. Ryzhikov.    Weak limits of powers of Chacon transformation. Preprint.  2012





\end{thebibliography}

\documentclass[a4paper,10pt]{article}
\def \Z {{\mathbf {Z}}}
\def \R {{\mathbf {R}}}
\def \N {{\mathbf {N}}}

\def \e {{\bf 1}}
\def\u{\bigsqcup}
\def\eps{\varepsilon}
\title{ Minimal Self-Joinings,  Bounded Constructions, \\ and  Weak Closure of Ergodic   Actions. }
\author{V.V. Ryzhikov}
\date{12.12.2012}

\begin{document}
\large
\maketitle
%\begin{abstract}{\Large 
For a weakly mixing bounded  rank-one  construction it is proved  the disjointness of its powers.  For non-rigid constructions we get  minimal self-joinings.  Examples of non-mixing rank one actions  with explicit weak closure are proposed. 
\section{Introduction} 
Consider rank-one measure-preserving transformations  of a Probability space $(X,\mu)$.
In connection with Thouvenot's question on minimal self-joinings
(MSJ) of mildly mixing rank-one transformations, for a class of transformations (mildly mixing bounded constructions) we prove the property MSJ. This class includes well-known classical and modified  Chacon's transformations (see\cite{JRS}).

The "springs" of our proof are similar to ones of \cite{JRS},  although  we do not consider generic points, and never  use Birkhoff's  ergodic theorem. Weak limits are our tools.  In operator terms
the method is following.  Let $P$ be an indecomposable Markov
operator commuting with a rank-one transformation $T$ ($P$ is  extreme point of  Markov centralizer of $T$).
If $P\neq T^i$ for all $i\in \Z$, then there is a sequence ${n_j}$
such that $T^{n_j}\to aP+(1-a)P'$, where  $a\geq \frac 1 2$,
$P'$ is some Markov operator.  In fact there is a sequence of
sets $Y_j$ such that $\mu(Y_j)\geq \frac 1 2$  and
$$\e_{Y_j}T^{n_j}\to aP, \ \ \  a\geq \frac 1 2.$$
Using non-rigidity of $T$   find a sequence $C_j\subset
Y_j$ for which
$$\e_{C_j}T^{n_j}\to cT^m P,\ \  \ \ c>0,  \ \ m\neq 0.$$
But from  $C_j\subset
Y_j$ we have  $\e_{C_j}T^{n_j}\to cP$ as well. This implies
$$P=T^mP.$$
If $T$ is mildly mixing, then  $T^m$ ($m\neq 0$) is ergodic,
hence, $P=\Theta$ -- the ortho-projector to constant functions in
$L_2(X,\mu)$.
We get the triviality of Markov centralizer. It is the set of all convex sums of the operators
$T^i$ and $\Theta$, consequently, $T$ has  MSJ.

In \cite{JRS} the authors used a similar trick: they 
proved for  a non-trivial self-joining $\nu$ of Chacon's map
the marginal invariance:
$$  \nu =(Id\times T)\nu, $$ 
so $  \nu =\mu\times \mu. $

We give   new  examples of non-mixing rank one actions  with explicit weak closure  that gives minimal self-joinings. 
An action  of such kind appeared in arXiv:1108.0568: for  double staircase transformation $T$  the  weak closure of its powers $Lim(T)$  is $$\{\Theta, 2^{-m}T^n+(1-2^{-m})\Theta \ : m\in\N,\ n\in \Z\}.$$ 
A non-simple description of power weak limits for  Chacon's transformation is given in \cite{JPRR}.

In this note we show that $Lim(T)$ for so-called stochastic Chacon's transformation (see \cite{R}) is (a.s.) 
$$\{\Theta, P^{m}{P^\ast}^{n}T^k \ : m,n\in\N,\ k\in \Z, m+n\neq 0,  P=aI+(1-a)T\}.$$ 

For bounded staircase flow $T_t$  we get
$$Lim(T_t)=
\{\Theta, T_a\prod_{i\in S} P_{m_i}\ :a\in \R,  S\subset \N, |S|<\infty,
 P_m=\int_0^mT_t dt\}.$$ 
\section{Rank-one constructions}
{ A rank-one construction} is determined by an integer $h_1$, a cut-sequence $r_j$ and  a spacer  sequence $\bar s_j$
$$ \bar s_j=(s_j(1), s_j(2),\dots, s_j(r_j-1),s_j(r_j)).$$  We recall its definition.
Let a transformation  $T$ is defined on the step $j$ as  a shift on a collection of disjoint sets (intervals)
$$E_j, TE_j T^2E_j,\dots, T^{h_j}E_j.$$
Cut $E_j$ into $r_j$ sets (subintervals)  of the same measure
$$E_j=E_j^1\u E_j^2\u  E_j^3\u\dots\u E_j^{r_j},$$  
then for all $i=1,2,\dots, r_j$ we  consider so-called columns
$$E_j^i, TE_j^i ,T^2 E_j^i,\dots, T^{h_j}E_j^i.$$
Adding $s_j(i)$ spacers over a $i$-th column we obtain  "column + spacers"
$$E_j^i, TE_j^i T^2 E_j^i,\dots, T^{h_j}E_j^i, T^{h_j+1}E_j^i, T^{h_j+2}E_j^i, \dots, T^{h_j+s_j(i)}E_j^i$$
(all intervals   are disjoint).
For all  $i<r_j$ we "stack"
$$TT^{h_j+s_j(i)}E_j^i = E_j^{i+1}.$$ 
Now one obtains a tower 
$$E_{j+1}, TE_{j+1} T^2 E_{j+1},\dots, T^{h_{j+1}}E_{j+1},$$
where 
 $$E_{j+1}= E^1_j,$$
$$T^{h_{j+1}}E_{j+1}=T^{h_j+s_j(r_j)}E_j^{r_j}, $$
$$h_{j+1}+1=(h_j+1)r_j +\sum_{i=1}^{r_j}s_j(i).$$

\bf Bounded   constructions. \rm  Following \cite {B}  consider  constructions  assuming  that   all $s_j(i)$, $r_j$ are bounded: $0\leq s_j(i)<s$,  $2<r_j<r$ .

(A reader can  slightly generalize the above notion, only requiring   for   all $j$ bounded "derivatives":
 $\max_{1\leq i< r_j} |s_j(i+1)-s_j(i)|<s, \ \ r_j<r.\ )$

If $s_j(r_j)=0$ for   all $j$, we get a "strongly bounded construction", which is  as a rule similar to  modified Chacon's transformation .
\\
\bf Examples. \rm

Odometers. For instance,  $r_j=5,$  $\bar s_j = (2,2,2,2, 0)$ for all $j$.

\rm Classic
Chacon's map:  $r_j=2$, 
  $\bar s_j = (0,1)$ for all $j$. It is not mixing, that is clear because of  the following  weak convergence
$$ 2T^{-h_j}\to \sum_{i=0}^{\infty}\left(\frac T 2 \right)^i. $$

 Modified Chacon's map (\cite{JRS}): $ r_j=3$, 
  $\bar s_j = (0,1, 0)$ for all $j$. Now 
$$ T^{-h_j}\to \frac {I+T} 2.$$

\section{ Joinings} \rm A {\it self-joining} (of order 2) is defined to be a $T\times T$-invariant
measure $\nu$ on $X\times X$ with the marginals  equal to  $\mu$:
$$\nu(A\times X)=\nu(X\times A)=\mu(A).$$
A joining $\nu$ is called ergodic if the dynamical system
$(T\times T, X\times X, \nu)$ is ergodic.
The measures $\Delta^i= (Id\times T^i)\Delta$ 
defined by the formula
$$
	\Delta^i(A\times B) = \mu(A\cap T^iB) 
$$ 
are referred to as {\it off-diagonals measures\/} (for $i\neq 0$). 
If $T$ is ergodic, then $\Delta^i$
are   ergodic self-joinings. 

We say that  $T$  has {\bf minimal self-joinings}
 of   order 2 (and we write $T\in MSJ(2)$)  if $T$  has no  ergodic
joinings except $\mu^{\otimes 2}=\mu\times\mu$ and $\Delta^i$.

The notion of MSJ (of all orders) has been introduced by D. Rudolph \cite{Ru} (see also \cite{T}).  In \cite{JRS} the authors  proved MSJ for 
modified Chacon's automorphism.  It is well-known fact that
MSJ(2)=MSJ for non-mixing maps (Glasner, Host, Rudolph, Ryzhikov).

The property of minimal self-joinings implies {\bf  mild mixing}. We recall that an automorphism $T$ is {\it mildly mixing} if  for any set $A$, $0<\mu(A)<1$,  
$$\limsup_j 
\mu(A\cap T^iA)  < \mu(A) .$$
An automorphism  $T$ is mildly mixing iff it has no rigid factors ( $S$ is rigid, if there is a sequence $m_j\to\infty$
such that $S^{k_j}\to Id$).  

 An automorphism $T$ is  {\it partially mixing}, if
 for for some $\alpha \in (0,1]$ and for all measurable sets $A,B$ 
$$\liminf_i 
\mu(A\cap T^iB)  \geq \alpha \,\mu(A)\mu(B).$$
% \vspace{2mm}
In \cite{KT} the authors proved minimal self-joinings for partially mixing
rank one transformations.
The property of partial mixing implies mildly
mixing. 

J.-P. Thouvenot conjectured  that all 
\bf bounded  mildly mixing  constructions have minimal self-joinings. \rm 
We give a positive answer (generalizing \cite{JRS} and some results from \cite{K}).
\vspace{3mm}

{\bf THEOREM.} \it  Let $T$ be a bounded non-rigid construction.  If it is totally ergodic (all non-zero powers are ergodic), then $T$ has minimal self-joinings.\rm
\vspace{3mm}

\section{Ergodic limits of off-diagonals are trivial}
\it Given  an  ergodic self-joining $\nu$ of a rank one transformation $T$, is   there  a sequence $n_j$ such that for all measurable $A,B$
$$\nu(A\times B) = \lim_{j\to\infty}   \mu(T^{n_j}A\cap B)      \ \ {\bf ?}$$
\rm This  question is due to J. King.

If such a sequence $n_j$ exists,  we call  $\nu$ 
a limit of off-diagonals and write
 $$\Delta_{T^{n_j}}\to \nu.$$

\bf LEMMA      \it Let $T$ be non-rigid totally ergodic bounded construction,  $m_j\to\infty$ and  $\Delta_{T^{m_j}}\to \nu.$  If $\nu$ is ergodic, then $\nu=\mu\times\mu$. \rm

Proof.  We find  $p=p(i)$  such that  $h_p\leq m_j <h_{p +1}$. Now we consider our construction on the step $p$.
Let's remark that $p=p(i)\to\infty$ and $rh_p>h_{p +1}$ for all (large)$p$.  The spacers  over the last column form 
the roof over $(p+1)$-th tower.  The transformation $T$ is non-rigid, hence the roof have to be (asymptotically) non-flat.  This implies the following:   $$T^{m_j}C_j\to cP,$$ and  for some Markov operator $Q$ commuting with $T$ we have 
$$P=(a_0I +a_1T +\dots+ a_{r-1}T^{r-1})Q,$$ where  at list two coefficients, say $a_n$, $a_{n+k}$, $k>0$, have to be both non-zero.   Let us rewrite this in joining terms:
$$\Delta_{T^{m_j}}\to \nu = \sum_{k=0}^{r-1}a_k(Id\times T^k)\eta.$$
 Thus, $\nu$ and $(Id\times T^k)\nu$ are not disjoint.
Assuming $\nu$ to be ergodic   we get
$$  \nu =(Id\times T^k)\nu. $$ 
From the ergodicity of $T^k$ it follows that
$$   \nu =\mu\times\mu $$ 
(in operator terms we say:  $T^kP=P \ \ implies\ \  P=\Theta.$)
Indeed,
$$   \nu(A\times B) =\int_{X\times X} \chi_A\otimes \left(\frac 1 N \sum_{d=1}^N T^{dk}\chi_B\right) d\nu =$$
$$ = \lim_N  \int_{X\times X}\chi_A\otimes \left(\frac 1 N \sum_{d=1}^N T^{dk}\chi_B\right) d\nu=
 \mu(B)\int_{X\times X}\chi_A\otimes 1 d\nu = \mu(A)\mu(B). $$

\section {Joinings as Local Limits}
 \rm Any ergodic self-joining  $\nu$ of a rank one transformation $T$  is  a partial limit of off-diagonals:   there is a sequence $n_j$ such that
$$\Delta_{T^{n_j}}\to \frac {1}{2} \nu  +\frac {1}{2} \nu'.$$
In fact (see  \cite{JRR}), for some 
$\delta \leq \frac {1}{2}$ there is a sequence of sets $Y_j$ in the form 
$$Y_j=\bigcup_{\delta h_j < k < h_j}T^kE_j$$  and a sequence $\{n_j\}$, $n_j \approx (1-\delta)h_j$,   such that 
$$\nu(A\times B) = \lim_{j\to\infty}  \frac {1}{\mu(Y_j)} \mu(T^{n_j}A\cap B\cap Y_j),$$
equivalently, 
$$\frac {1}{\mu(Y_j)} \e_{Y_j}T^{n_j}\ \to\  P.$$
Such limits we call local.
The following lemma shows that sometimes certain local limits become global.
 
\bf LEMMA 1. \it  Let $\e_{Y_j}T^{n_j}\to (1-\delta)P$,  $h_j\leq n_j<h_{j+1}$.
We represent  
$$n_j = q h_{j} +s_j(1)+s_j(2) + \dots + s_j(q) + m_j, \ \  0\leq m_j <h_j. $$

  If $\frac {m_j}{ h_{j}}\to 0$ (or $ \frac {h_{j}-m_j}{ h_{j}}\to 0$),  then 
$T^{m_j}\to P$ (or $T^{m_j-h_j}\to P$). \rm

Proof.  Let $C_j^1$ denote  the first column of $j$-tower.
We see that  $$\mu(C_j^1)>\frac 1 r, \ \ \
T^{n_j}C_j^1 \subset Y_j, \ \ \ \mu(T^{n_j}C^1_j\Delta C^{q+1}_j)\to 0.$$
We get
$$\nu(A\times B) = \lim_{j\to\infty}  \frac {1}{\mu(Y_j)} \mu(T^{n_j}A\cap B\cap Y_j)=$$

$$= \lim_{j\to\infty}  \frac {1}{\mu(C^1_j)} \mu(T^{n_j}A\cap
T^{n_j}C^1_j\cap B)=$$
$$= \lim_{j\to\infty}  \frac {1}{\mu(C^1_j)} \mu(T^{m_j}A\cap
C^{q+1}_j\cap B)
=\lim_{j\to\infty}  \mu(T^{m_j}A
\cap B).$$

\section{On local "breaking"  of non-trivial  self-joinings }

    Assume  $T$ satisfies the conditions of Theorem 1.  Let's find 
sequences of sets  $C_j', C_j''\subset Y_j$ such that 

\ \ \ \ \ \ \  \ \ \ \  $\mu(C_j')\approx \mu(C_j'')\to c>0$,\ \ $  \e_{ C_j'}{}^{{}_\circ} T^{n_j}\to cP$,\ \
$  \e_{ C_j''}{}^{{}_\circ} T^{n_j}\to cP$,
\\ and in addition  for fixed  $ m>0$ and some  sequence $m_j$,  $|m_j|<h_j$,   let
$$  \e_{ C_j'}{}^{{}_\circ} T^{n_j}\approx_w \e_{ C_j'}{}^{{}_\circ}  T^{m_j},\eqno (*) $$
$$   \e_{ C_j''}{}^{{}_\circ} T^{n_j}\approx_w  \e_{ C_j''}{}^{{}_\circ}T^{m_j +m},\eqno (**)$$ 
$$  \e_{  C_j'}{}^{{}_\circ} T^{m_j}\approx_w \e_{ C_j''}{}^{{}_{\circ }} T^{m_j}. \eqno (***)$$
If such $C_j', C_j''$ are found,  then
$$cP\approx_w \e_{ C_j''}{}^{{}_\circ}T^{m_j +m} \approx_w
\e_{ C_j'}{}^{{}_\circ}T^{m_j +m}\approx_w cT^mP, \ \
P=T^mP, \ \ P=\Theta.$$

Repeat the above in joining terms : 
 for all measurable $A,B$
$$\nu(A\times B) =
 \lim_{j\to\infty}   \mu(T^{n_j}A\cap B\cap C''_j)/{\mu(C''_j)}\\
=_{(**)}  \lim_{j\to\infty}  \mu(T^{m_j+m}A\cap B\cap C''_j)/{\mu(C''_j)}=_{(***)}$$
$$  \lim_{j\to\infty}   \mu(T^{m_j}T^m A\cap B\cap C'_j)/{\mu(C'_j)} =_{(*)}
\lim_{j\to\infty}   \mu(T^{n_j}T^m A\cap B\cap C'_j)/{\mu(C'_j)}=\nu(T^m A\times B).$$
Thus, 
$  \nu =(T^m\times Id)\nu,   \ \ \  \nu=\mu\times\mu. $

\bf How to find  $C_j', C_j''$?
\rm
Consider  minimal  $i$ for which $s_j(i)\neq s_j(i+1)$.  Without loss of generality  suppose $i<0.4 r_j$.

{\normalsize In oder to have this, instead of the sequence
$\{j\}$ (of steps)  we may consider  the subsequence  $\{Nj\}$,    where $N$ is fixed.  Now   $$\tilde r_{j}:= \prod _{k=(j-1)N+1}^{jN}r_{k}.$$  
 Let's look at Chacon's map  spacer sequence  $\bar s_j = (0,1, 0)$.
Setting $j:=2j$, we get new spacer sequence  
$\bar{\tilde s}_j= (0,1,0,0,1,1,0,1,1). $
Generally,  given construction $T$,  if for all $N$  we cannot find $i$ s.th.   $s_j(i)\neq s_j(i+1)$, then we see  flat  roofs  of towers,   hence,  $T$ have to be  rigid. It is not our case. Thus, we get 
for some $N$  for all (new) $j$  a desired integer $i=i(j)$  satisfied  $s_j(i)\neq s_j(i+1)$. If $N$ is sufficiently large, we find $i<0.4r_j$. }

If  the image under $T^{n_j}$ of the left half  of $E_j$ ( and the left part of spacers)  is
located far away from the top and bottom of the tower,
then we easy find   $C_j', C_j''$ as below.  Let's remark that in the case of 
the  flat image  at the top of the tower, we are looking for $C_j', C_j''$ at the bottom  (and vice versa, respectively).

 {%\small
\begin{picture}(0, 600)
%\put(200,600){\it \Large The sets $C'_j$, $C''_j$ up }
\put(42,0){\vector(2,3){230}}
\put(0,0){  \line(1, 0){390}}
%\put(-4,600){  \line(1, 0){390}}
\put(16,600){  \line(1, 0){40}}
\put(16,610){  \line(1, 0){20}}
{\Large
\put(65,605){spacers}}

{\small
\put(365,605){spacers}}
%\put(0,0){  \line(0, 1){600}}
%\put(400,0){  \line(0,1){600}}

\multiput(0,0)(20,0){11}{
\multiput(0,0)(0,10){60}%
{${.}$}}
\multiput(215,0)(20,0){10}{
\multiput(0,0)(0,10){60}%
{${.}$}}

%\multiput(400,0)(0,10){60}%
%{${\bf .}$}
\multiput(0,0)(0,40){15}
{  \line(1,0){200}}

\multiput(210,20)(0,40){8}
{  \line(1,0){180}}
\put(210,299){  \line(1, 0){180}}

%\put(250,301){  \line(0, 1){290}}

\put(40,0){\vector(2,3){230}}
\multiput(210,580)(0,-40){7}
{  \line(1,0){40}}
\multiput(270,590)(0,-40){7}
{  \line(1,0){20}}
\multiput(270,590)(0,-40){6}
{  \put(8,-25){{\Large \bf C$''_j$}}}
\multiput(250,560)(0,-40){6}
{  \put(7,15){{\Large \bf C$'_j$}}}
\multiput(250,560)(0,-40){6}
{  \line(1,0){20}}
\multiput(290,580)(0,-40){7}
{  \line(1,0){100}}

\put(0,561){  \line(1, 0){200}}
%\put(0,558){  \line(1, 0){200}}

%\put(210,421){  \line(1, 0){40}}
%\put(210,219){  \line(1, 0){40}}
%\put(290,221){  \line(1, 0){100}}
%\put(290,219){  \line(1, 0){100}}

%\put(250,201){  \line(1, 0){20}}
%\put(250,199){  \line(1, 0){20}}

%\put(270,211){  \line(1, 0){20}}
%\put(270,210){  \line(1, 0){20}}
%\put(270,209){  \line(1, 0){20}}
%\put(130,100){  \li,ne(1,-2){50}}
{\Large
\put(135,5){${\bf  E_j}$}
		\put(115,45){${\bf  T^4E_j}$}
\put(115,85){${\bf T^8 E_j}$}
}

{\Large
\put(330,345){${\bf T^{n_j} E_j}$}
\put(395,350){level $ \ {\bf {m_j}}$}
		\put(320,385){${\bf T^{n_j+4} E_j}$}
\put(320,425){${\bf  T^{n_j+8}E_j}$}

}

{\Large
\put(115,565){${\bf  T^{h_j -4}E_j}$}
		\put(315,305){${\bf  T^{n_j+h_j -4}E_j}$}}
%\put(150,-12){$r_j=19$} }
\end{picture}
%  \begin{center}Fig. 1\end{center}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\begin{picture}(0, 630)
\put(0,-20){\it \large  If the  image  at the top of the tower is flat, we are looking for $C_j', C_j''$ at the bottom  }
%\put(22,0){\vector(1,1){170}}
\put(80,560){\vector(2,-3){222}}
\put(0,0){  \line(1, 0){390}}
%\put(-4,600){  \line(1, 0){390}}
\put(16,610){  \line(1, 0){40}}
\put(16,620){  \line(1, 0){20}}
{\Large
\put(65,605){spacers}}

{\small
\put(365,605){spacers}}
%\put(0,0){  \line(0, 1){600}}
%\put(400,0){  \line(0,1){600}}

\multiput(0,0)(10,0){40}{
\multiput(0,0)(0,10){60}%
{${. }$}}

%\multiput(400,0)(0,10){60}%
%{${\bf .}$}
\multiput(0,0)(0,40){15}
{  \line(1,0){200}}

\multiput(210,260)(0,40){9}
{  \line(1,0){180}}

%\multiput(270,590)(0,-40){7}
%{  \line(1,0){20}\put(-15,-25){{\Large \bf C$''_j$}}}
%\multiput(250,560)(0,-40){6}
%{  \line(1,0){20}\put(-15,-25){{\Large \bf C$'_j$}}}

\multiput(210,260)(0,-40){7}
{  \line(1,0){40}}
\multiput(270,250)(0,-40){7}
{  \line(1,0){20}}
\multiput(250,240)(0,-40){7}
{  \line(1,0){20}}
\multiput(250,240)(0,-40){6}
{  \put(6,-25){{\Large \bf C$'_j$}}}
\multiput(250,240)(0,-40){6}
{  \put(26,-15){{\Large \bf C$''_j$}}}
\multiput(290,260)(0,-40){7}
{  \line(1,0){100}}

\put(0,559){  \line(1, 0){200}}
\put(0,558){  \line(1, 0){200}}

\put(210,221){  \line(1, 0){40}}
\put(210,219){  \line(1, 0){40}}
\put(290,221){  \line(1, 0){100}}
\put(290,219){  \line(1, 0){100}}

\put(250,201){  \line(1, 0){20}}
\put(250,199){  \line(1, 0){20}}

\put(270,211){  \line(1, 0){20}}
\put(270,210){  \line(1, 0){20}}
\put(270,209){  \line(1, 0){20}}
%\put(130,100){  \li,ne(1,-2){50}}
{\Large
\put(35,5){${\bf  E_j}$}
		\put(15,45){${\bf  T^4E_j}$}
\put(15,85){${\bf T^8 E_j}$}
}

{\large
\put(255,265){$ \ {\bf T^{n_j} E_j}$}

		\put(245,305){${\bf T^{n_j+4} E_j}$}\put(245,345){${\bf  T^{n_j+8}E_j}$}

}

{\Large
\put(15,565){${\bf  T^{h_j -4}E_j}$}
		\put(315,225){${\bf  T^{n_j+h_j -4}E_j}$}
%\put(150,-12){$r_j=19$}
 }
\put(15,250){\large level $ \ {\bf {h_j+m_j, \ m_j< 0}}$}
\end{picture}
\\ \\ \\ \\ \\
If  the mentioned image of  $E_j$  is close to the top, or to the  bottom,  we use Lemma 1. This finish the proof of Theorem 1.
\newpage

}
\Large
\section{ Bounded constructions and disjointness of powers}
In  \cite{B} J. Bourgain  proved 
 	that bounded constructions  satisfied the Moebius orthogonality property.  This property is a consequence of MSJ  (in fact, of the disjointness  of $T^q$ and $T^p$, $q\neq p$). 

\bf THEOREM 2. \it   If a bounded construction $T$ is weakly mixing and $q\neq p$,  then 
$T^q$ and $T^p$ are disjoint.\rm

Proof.  
 If  roofs over last columns are asymptotically  non-flat, a weakly mixing construction is not rigid, hence,  it has MSJ (Theorem 1).
If   a flatness appears, this means that  our spacer sequence is following: for some sequence of integer intervals  $[\alpha_t, \beta_t]$  $( \beta_t-\alpha_t) \to\infty$

$$ \bar s_j = (s_t, s_t,\dots, s_t, s_t, 0), \ \  \forall  j\in [\alpha_t, \beta_t]      
 $$ 
( "0"-s are necessary in this situation).
From the weak mixing property,  from time to time we see  a break in  the mentioned  behavior (our construction is not an odometer!).
 Thus, infinitely many times  we meet  non-flat roofs over last columns.
Given $\eps>0$, it is not hard  to get for  all $k<\eps^{-1}$  the following weak limits: 
$$T^{kn_j}\to  (1-k\eps) I + k\eps P,$$
where $P=\sum_{i\geq 0}^{\infty} a_i T^i,$   
$\sum_{i\geq 0}^{\infty} a_i =1$, and $a_m> 0$   for some $m>0$.  

If $$T^qJ=JT^p$$ for an operator $J:L_2(\mu)\to L_2(\mu)$, then
$$((1-q\eps) I + q\eps P)J=J((1-p\eps) I + p\eps P).\eqno (\ast)$$

Assume  $J$ to be an indecomposable Markov operator (indecomposable
in the convex set of Markov operators intertwining $T^q$ and $T^p$, or,
in other words,  the corresponding joining is ergodic). All operators $T^nJ,  JT^m$ are indecomposable as well.
Let  $q<p$, from $(\ast)$ it follows $J=JT^{m'} $  for some $m'>0$ (again a  marginal invariance of a joining!).    The ergodicity of 
$T^{m'}$  implies $J=\Theta$.  Thus,  $T^q$ and $T^p$ are disjoint.
 
\large
%%%%%%%%%%%%%%%%%%%%%%%%\put(205,305){$the %\ image \ {\bf T^{n_j} E_j, {\it \ at\ the\ level} \ \  T^{m_j}E_j}$}
\section{Explicit Weak Closure of Actions}
\bf Stochastic Chacon's map. \rm 
Fixing $h_1$,  cut-numbers $r_j\to\infty$,   consider  the set 
of all constructions with spacer sequences
  $s_j(i)\in \{ 0, 1\}$. Let's equip this ensemble
  by Bernoulli measure of type 
$ (a, 1-a)$.
I.e., in a random fashion, with probability $1-a$
 we stack
one spacer  over a column.  Denote   $P=aI+(1-a)T^{-1}$.

\bf THEOREM 3. \it
For  stochastic Chacon's maps $T$  almost surely 
$$ Lim(T)=\{\Theta, P^{m}{P^\ast}^{n}T^k \ : m,n\in\N,\ k\in \Z, m+n\neq 0\}.$$ 

\bf COROLLARY.  \it $ Lim(T)=  \{ \Theta,   T^kP^n : n=0,1,2 ....   n >0, k \in \Z\}$ as $a=0.5$.
\rm

\bf LEMMA. \it 
Let  $T^{n_j} \to Q$,                   $ n_j >0$. 
Represent
$n_j =  q_j h_{p(j)}   +      m_j$ , 
    where    $h_{p(j)}\leq  n_j < h_{p(j)+1}$.    $ 0\leq  m_j \leq  h_{p(j)}$,     $q_j \leq  r_{p(j)}$.
   If $q_j\to\infty $  and  $(r_{p(j)} -q_j)\to  \infty$,  then $Q=\Theta$.
\\
\rm
The proof of this lemma uses the following facts:
For almost all T
for any $q>0$
    $$T^{qh_j} \to  P^q $$ 
and
$$P^q \to \Theta,  \ \ as \ \ q\to \infty,$$ moreover,
  $T^n P^q \to \Theta$  uniformly with respect to $n$.

Assume that $q_j$ (or $r_p(j)- q_j$) to be  bounded.
If   $\eps h_{p(j)}< | m_j |<(1-\eps)h_{p(j)}$, then $Q=\Theta$.  
Let $$\frac {m_j}   {h_p(j)}  \to   0.$$
For all $i$, $0<i < r_{p(j)}-q+1$
$$T^{m_j  -s_j(i)-s_j(i+1)-\dots- s_j(i+q-1)} \approx_w Q,$$
$$T^{n_j}\approx_w \frac 1 {r_{p(j)}} \sum_{i=1}^{r_{p(j)}-q} T^{ -s_j(i)-s_j(i+1)-\dots- s_j(i+q-1)}T^{m_j } \approx_w P^q T^{m_j}.$$

Thus,  $Q=       P^q   \lim_j   T^{m_j}$, then we get 
 $$ Q= P^k {P^*}^n Q^{''}$$
and so on. Note that ${P^*}$ appears in case  of  
$\frac {m_j}   {h_p(j)}  \to   1.$
Iterating, if we cannot stop  by   $ Q^{'''\dots '''}=T^m$, then  $P=\Theta$  ($P^k {P^*}^nT^m \to \Theta$ as $k+n\to \infty$  uniformly with respect to $m$).

\vspace{5mm}
\bf  Bounded staircase flow $T_t$. \rm
Fix $h_1\in R^+$ and cut-sequence $r_j\to\infty$. Define  spacers  by
  $s_j(i)= \frac  {i-1} {r_j}$  ($1\leq i\leq r_j$).

We have 
$$T_{mh_j}\to
 P_m:=\int_{-m}^0T_t dt=T_{-m} P^*_m.$$

\bf THEOREM. \it  For bounded staircase flow $T_t$  
$$Lim(T_t)=
\{\Theta, T_a\prod_{m\in M} P_{m}\ :a\in \R, \ M\subset \N, \ |M|<\infty
 \}.$$ 
\rm

\bf LEMMA. \it 
Let  $T^{t_j} \to Q$,                   $h_{p(j)}\leq  t_j < h_{p(j)+1}$.   
Represent
$t_j =  q_j h_{p(j)}   +      m_j$ , 
    where  $ 0\leq  m_j \leq  h_{p(j)}$,     $q_j \leq  r_{p(j)}$.
   If $q_j\to\infty $  and  $(r_{p(j)} -q_j)\to  \infty$,  then $Q=\Theta$.
\rm

The  proof of the  lemma  is an exercise.

Let  $T_{t_j}\to Q\neq \Theta$,  then Lemma
asserts that there is $q$  such that $t_j =  q h_{p(j)}   +      m_j$.
Again we have 
$$\frac {m_j}   {h_{p(j)}}  \to   0\ \ or \ \ \frac {h_{p(j)}-m_j}   {h_{p(j)}}  \to 1,$$
and
$$ Q= P_q \lim T_{m_j} \ \ \left( \ \ or \ \ Q= {P}_{q+1} \lim T_{m_j-h_{p(j)}}\right).$$
Iterating, taking into account  the fact that 
$T_m\prod_{q\in M} P_{q} \to\Theta $ uniformly with respect to $m\in \R$ as $|M|\to \infty$, we stop as $Q= T_a\prod_{m\in M} P_{m}$.

\begin{picture}(650, 490) 
\multiput(0,0)(8,2){25}%
{  \line(0, 1){400}   }
\multiput(0,400)(8,2){24}%%%%%%%%%%%%
{  \line(1, 0){10}   }
\multiput(0,-50)(8,0){25}%
{  \line(0, 1){200}   }
\put(4,-50){\line(1,0){200} }
%\put(2,-49){\line(1,0){198} }
%\put(5,-9){1}
%\put(13,-9){2}\put(21,-9){3}
\multiput(16,300)(8,-4){22}% %%%%%%%%%%%%%
{  \line(1, 0){10}   }
\put(4,215){\line(1,-5){10}}
\put(4,216){\line(1,-5){10}}

\put(12,265){\line(1,-5){10}}
\put(12,266){\line(1,-5){10}}
\put(190,445){\line(1,5){10}}
\put(189,445){\line(1,5){10}}
\put(197,444){\line(0,1){50}}
%\multiput(112,492)(8,-24){10}%
%{  \line(1, 0){10}   }
%\put(38,290){1}
%\put(46,258){2}\put(54,226){3}
%{\large 
%\put(228,140){What is a delay?  If $T^m E^i_j$ is situated in  $T^tE_j$, then }
%\put(228,125){$T^m E^{i+1}_j$ will be  in  $T^{t-d}E_j$. }
%\put(224,110){ This $d$ is an ``upper'' relative  delay}
\put(28,260){  $ T_{t_j}E_j$}
\put(210,300){$T_{m_j}E_j$}
\put(210,395){$T_{h_j}E_j$}
\put(210,-52){$E_j$}
\put(0,-62){$q=2$}
%\put(0,-65){  $E_j^1$, $E_j^2$,$E_j^3$,$\dots$}
%\put(39,15){ delay = 5 }\put(119,145){ delay = 4 }\put(140,445){ delay = 3 }
%\put(250,220){$m =\sum_{i=1}^{d_j} ( h_{j}+i-1) + t_j$
%  } 

\put(3,302){\line(1,0){200} }   \put(3,400){\line(1,0){200} } 
%\put(35,301){\line(1,0){8} }  

\end{picture}  
\Large
\newpage
\section{ Related Problems}

\
1,2.  Thouvenot's questions:   Do    mildly mixing rank-one transformations
possess the properties $PID$? $MSJ(2)$?
\vspace{4mm} \\
3. Is
 $Rank(T^n)$ for  mildly mixing rank-one transformation $T$ equal to $n$ ($n>1$)?   
\vspace{3mm} \\
4. King's  question:  will  any ergodic self-joining of a rank-one transformation be  a limit of off-diagonals measures?
\vspace{3mm} \\
5. Is it true that for any mildly mixing bounded construction $T$ its
symmetric powers $T^{\odot n}$ have simple spectrum? (see \cite{R})
\vspace{4mm} \\
6,7. Let $T$ be  Chacon's map.  Will  the product
$T\otimes T^{2}\otimes T^{3}
\otimes T^{4}\dots $ be of
 simple spectrum  for Chacon's map ? for a mildly mixing  bounded construction? 
\vspace{2mm} \\
8. Consider a rank-one flow called rigid  staircase flow. 
Fix $h_1\in R^+$ and a cut-sequence $r_j\to\infty$. Define  spacers  by
  $$s_j(i)= \frac  {i-1} {jr_j}, \ \   i\leq r_j.$$ Will the corresponding rank-one flow be  simple ?  (see  \cite{JR1} or \cite{T} for definitions  of simplicity. An example of a rigid, simple transformation is given in  \cite{JR}.)  
\vspace{2mm}
\\
9. Let $T$, $T'$  be rank-one constructions possessing  the same  spacer sequences, and $h_1\neq h_1'$ for them.  Are  they disjoint ? 
(It is true for mixing $T$, $T'$, see  	arXiv:1109.0671.)
\vspace{2mm}
\\
10. Are there mildly mixing  (bounded) constructions $T$, $T'$ such that $T'$ is  spectrally isomorphic to $T$  but not isomorphic (as measure-preserving map)  to $T$ and $T^{-1}$?  For mixing rank-one flows  this is
possible:  the flows $T_t$ and $T_{\alpha t }$ ($\alpha >1$)  from  
 arXiv:1002.2808  have the same  spectrum (A. A. Prikhodko), but they are disjoint (see  	arXiv:1109.0671).
\vspace{2mm}
\\
11. For rank-one transformations $MSJ(2)=MSJ$.   Recall that MSJ can be defined as  $MSJ(2)\cap PID$, where the  property $PID$ (pairwise independence
determines  the  global independence)  means  
that any pairwise independent self-joining  have to be trivial (a product measure). This property has been introduced in   \cite{JR1} (see also
\cite{T}).   

  Let $\nu$ be a 3-fold self-joining of weakly mixing rank-one map $T$, and 
for all measurable $A,B\subset X$
$$\nu(A\times B\times X)=\nu(X\times A\times B)=\nu(A\times X\times B)=
\mu(A)\mu(B).$$
Is it true that $\nu= \mu\times \mu\times \mu \ ?$
\newpage
12. Does the condition  $T^{m_j}, T^{n_j},T^{m_j-n_j}\to\Theta$  implies   
$$\mu(A \cap T^{m_j}B \cap  T^{n_j}C)\to \mu(A)
\mu(B)\mu(C)$$
 for a mildly mixing rank-one map $T$?

%\normalsize

 E-mail: vryzh@mail.ru
\end{document}